# Threshold-less and Flexibly Tunable Frequency Comb via Floquet Engineering

### v2

@@ Abstract
Frequency combs have revolutionized communication, metrology and spectroscopy. Numerous efforts have been dedicated to developing integrated combs, predominantly relying on Pockels or Kerr mechanisms. In this work, we propose and demonstrate a new type of frequency comb—Floquet cavity frequency comb—that does not rely on intrinsic material non-linearity. By periodically modulating the resonance frequency of a cavity, a Floquet cavity with multiple equally spaced frequency components is created. The pump tone interacts with the pre-modulated cavity, generating the output frequency comb. This approach offers a flexible tuning range and operates in a threshold-less manner, obviating the need to overcome nonlinear initiation threshold. We implement this on a microwave cavity optomechanical system on-chip. Compared to Kerr optomechanical combs, this approach efficiently generates comb with pump signal far from the cavity’s intrinsic frequency, and the power required for detection is reduced by approximately a factor of $10^{6}$, providing a promising platform for frequency comb generation.
@@ Introduction
Figure 1: Floquet cavity and frequency comb. a Examples of cavities with periodically modulated resonance frequencies. b Model of the Floquet cavity with the modulation frequency $\Omega$ and strength $A$. c Time evolution of the intracavity field amplitude in the Floquet cavity obtained via numerical simulation. For demonstration purposes, cavity dissipation is neglected and $A\approx 25\Omega$. d The relative strength in the frequency domain of the Floquet cavity modes $\omega_{0}+m\Omega$ (where $m$ is the index of the Floquet states). The red dotted line marks the intrinsic frequency $\omega_{0}$. e Schematics of the spectral density of Floquet cavity quasi-energy with increasing modulation strength $A$ (where $n$ is the Fock basis). f The schematic diagram of the generation of Floquet cavity frequency comb. By injecting the pump tone into the pre-modulated cavity, photons are emitted across multiple modes, resulting in a frequency comb.
@@
Frequency combs, composed of discrete, equally spaced frequencies [1, 2], have contributed to advancements in optical communication [3, 4], precision metrology [5, 6], spectroscopy [7, 8], and atomic clock [9, 10, 11]. The compact on-chip frequency combs (microcombs) include microresonator-based combs [12, 13, 14] and electro-optic (EO) combs [15, 16, 17, 18, 19]. Microresonator combs relying on Kerr nonlinearity have been successfully demonstrated in silica [12], silicon nitride [20], silicon carbide [21], diamond [22], and lithium niobate (LN) [23]. Integrated EO combs utilize the Pockels effect, which have also recently been realized in thin-film LN devices [15, 16, 17] and lithium tantalate [19]. Despite these remarkable advances, EO and microresonator combs are facing limitations of their generation efficiency [15, 24, 25] or tuning precision [26, 20]. Therefore, emerging paradigms for frequency comb generation that do not rely on intrinsic medium nonlinearity are highly desired [27, 28, 29], as they offer promising avenues to overcome the aforementioned limitations in applications such as astrocombs [30, 31] and molecular fingerprinting [7, 32]. The frequency combs based on engineered nonlinearity could also offer a scalable and low-cost manner for emerging fields such as integrated photonic data links [33], quantum networks [34], and multiphoton entanglement [35, 36, 37]. Notably, Floquet engineering [38], which involves the periodic time-dependent modulation, has the capability to control and generate new multi-mode sidebands by tuning energy spectra [39, 40, 41], as well as to induce interactions between modes [42, 43]. These Floquet sidebands demonstrate that Floquet engineering is a promising although underexplored approach for frequency comb generation, without necessitating Kerr nonlinearity or Pockels effect.
@@
In this work, we propose and demonstrate a new type of frequency comb, termed the Floquet cavity frequency comb. Specifically, by periodically modulating the resonance frequency of a cavity, we create a Floquet cavity with multiple equally spaced frequency components. Then, a single-frequency pump tone is added, interacting with this pre-modulated multi-mode structure to generate the output frequency comb. Additionally, the absolute positions of the comb teeth are referenced to the frequency of the pump tone, with the intensity approximately symmetrically distributed around the cavity’s intrinsic frequency. Furthermore, in contrast to Kerr or EO combs, which require precise frequency tuning to match the free spectral range, our approach is flexibly tunable, with the spacing and number of comb teeth determined by the modulation source. The frequency components of the Floquet cavity are predefined, enabling the frequency comb to exceptionally operate without the need to overcome a threshold condition.
@@
Based on the proposed theoretical framework, such a frequency comb can be realized across a wide range of physical systems. Here, we demonstrate its implementation for modulating the resonance frequency of a superconducting microwave cavity via a mechanically compliant capacitor. To achieve sinusoidal modulation, we additionally employ an auxiliary cavity to establish a conventional optomechanical coupling, driving the mechanical oscillator from thermal state into self-induced oscillation [44]. In contrast to Kerr optomechanical combs [45, 46, 47], which rely on the nonlinear coupling between phonons and photons [48, 49, 50], our approach achieves a significantly extended tuning range for the pump signal, including frequencies that are substantially red-detuned from the cavity’s natural resonance. Furthermore, this approach reduces the input power required for detection by a factor of approximately $10^{6}$, which is limited primarily by the detection efficiency. This positions Floquet cavities as a promising platform for frequency comb generation.
@@ Results
We introduce a general Floquet cavity, defined as a cavity whose resonance frequency $\omega_{\mathrm{c}}$ is periodically modulated in time. It can be mapped to counterparts in well-established physical systems, including Fabry–Pérot resonators of movable mirror or microwave circuits with adjustable capacitance or inductance, as illustrated in Fig. 1a. The resonance modes of these Floquet cavities transform from a single-frequency state to a superposition of multiple frequencies (see Supplementary Note I.A). The Hamiltonian of the Floquet cavity is given by $H_{\mathrm{Floquet}}=\mathrm{\hbar}\left(\omega_{\mathrm{0}}+A\cos(\Omega t+\phi)\right)\hat{a}^{\dagger}\hat{a}$, where $\omega_{\mathrm{0}}$ is the intrinsic frequency, $A$ is the modulation strength, $\Omega$ is the modulation frequency, and we drop an irrelevant phase $\phi$, as illustrated in Fig. 1b.
In terms of Floquet theory, such periodic modulation of the cavity frequency effectively creates “Floquet sidebands” around the cavity’s intrinsic frequency. The quasi-energy of this Floquet cavity satisfies $\epsilon_{nm}/\hbar=n\omega_{\mathrm{0}}+m\Omega$, with the eigen-frequencies $\omega_{\mathrm{0}},\omega_{\mathrm{0}}\pm\Omega,\omega_{\mathrm{0}}\pm 2\Omega,\dots$(see Methods), where integer $n$ is the Fock basis of bare cavity and $m$ is the Floquet index. We present a numerical example illustrating the time evolution of the intracavity field in the Floquet cavity in Fig. 1c. A set of equally spaced components corresponding to the intracavity field in the Fourier domain is shown in Fig. 1d. The relative strength of the Floquet modes is determined by $J_{\mathrm{m}}(\beta)$, the Bessel function of the first kind with order $m$ (see Supplementary Note I.B and I.C). Here we define the dimensionless modulation coefficient $\beta=\frac{A}{\Omega}$. With a sufficiently large modulation strength, a substantial number of detectable Floquet modes can be produced.
Remarkably, since these modes are induced by Floquet engineering, they are not independent (as in modes in spatially separate cavities) but are correlated. The modulation strength $A/2$ determines the coupling strength between the different Floquet modes, thus forming what is referred to as a giant-mode cavity. A stronger modulation strength clearly enhances the generation of additional sidebands with appreciable intensity, as shown in Fig. 1e. Furthermore, this coupling enables energy input into the cavity to be transferred to different sidebands. Consequently, any input pump tone, even in the few-photon regime, is emitted across multiple sideband modes, forming a frequency comb with spacing $\Omega$, as illustrated in Fig. 1f.
Considering a input pump tone at the frequency $\omega_{\mathrm{in}}$, from the equation of motion, we obtain that the complex amplitude in the Floquet cavity satisfies (Supplementary Note II.A):
@@
×$\times\sum_{m=-\infty}^{\infty}\sum_{l=-\infty}^{\infty}\frac{J_{l}(\beta)J_{m-l}(-\beta)}{il\Omega+\kappa/2-i\Delta}e^{im\Omega t},$
where $\Delta=\omega_{\text{in}}-\omega_{0}$ arises from the rotation of the frame, $\kappa$ represents the total dissipation of the cavity, $\kappa_{\mathrm{e}}$ denotes the external dissipation due to the coupling between the cavity and the external input/output, and $a_{\mathrm{in}}$ represents the externally injected pump field.
For a Floquet cavity, the modulation coefficient $\beta$ is predetermined and independent of $a_{\mathrm{in}}$, enabling threshold-less frequency comb operation, as indicated by Eq. (1). In other terms, the pump tone does not induce nonlinear effects but instead serves as a probe for a system with a prescribed time-varying response function. Furthermore, we found that the pump tone exhibits highly flexible tuning capabilities within the approximate range $-2\beta\Omega\lesssim\Delta\lesssim 2\beta\Omega$, as detailed in the Supplementary Note II.A. When $\Delta\neq 0$, the pump tone alters the relative intensities of the frequency comb components to some extent. In contrast, in Kerr optomechanical combs, the generation of the frequency comb by optomechanical non-linearity depends on the pump tone itself, meaning $\beta$ is influenced by both $a_{\mathrm{in}}$ and $\Delta$. Consequently, $a_{\mathrm{in}}$ must reach a sufficient intensity to overcome an initiation threshold. Additionally, $\Delta$ needs to be constrained within a range that can effectively excite the oscillator (For $\kappa<\Omega$, $\Delta\approx\Omega$) in Kerr optomechanical combs.
Figure 2: Construct the Floquet cavity in optomechanical systems. a Micrographs of the device, showing the entire chip (top) and zoomed-in drums (bottom). We select the larger drum (left) for the demonstration, while the other one is ignored. b Equivalent circuit model: straight arrows indicate the corresponding applied signals (blue for cavity-1, purple for cavity-2). c Cavity-1 serves as the auxiliary driving cavity to drive the mechanical oscillator into self-induced oscillation. Cavity-2 is modulated by the oscillator to construct the Floquet giant-mode cavity. d Mechanical oscillator spectrum transitioning from thermal state to self-induced oscillation regime; the red curve marks the spectrum at the critical drive power $P_{\mathrm{dc}}$. e The reflection spectrum of the cavity-2 measured by S21 as a function of increasing drive power. Colored dots: measurements; black lines: fits based on the reflection coefficient $R$. $X_{\mathrm{0}}$ is incorporated into $\omega_{\mathrm{c2}}$. The inset marks the driving frequency schematics.
@@
We demonstrate Floquet cavity frequency comb in a superconducting microwave circuit, as illustrated in Fig. 2a. The resonance frequency of the cavity is modulated by a displacement-dependent capacitor. The sample used in this study is similar to that reported in our previous work [51, 52] (detailed fabrication in Supplementary Note III.A). Our device consists of two aluminum drumhead mechanical oscillators [53, 54] optomechanically coupled to two microwave cavities, as schematically shown in Fig. 2b. Here, we ignore one of the mechanical oscillators. The Hamiltonian of such a system can be expressed as
@@
where $\omega_{\mathrm{c},j}$ represents the $j$-th cavity’s intrinsic frequency. $\omega_{\mathrm{m}}$ represents the mechanical resonant frequency, and $a_{j}$ ($a_{j}^{\dagger}$) and $b$ ($b^{\dagger}$) represent the annihilation (creation) operators of the cavities and mechanical oscillator, respectively. $\hat{x}$ is the displacement of the mechanical oscillator, and $x_{\mathrm{zpf}}$ represents the zero-point fluctuation of the mechanical oscillator’s displacement. Besides, the coefficient $g_{\mathrm{0},j}$ represents the single photon coupling strength between the $j$-th cavity and mechanical oscillator.
Our entire device operates in a dilution refrigerator at a temperature of 10 mK (detailed setup in Supplementary Note III.B). The first cavity’s (cavity-1) intrinsic frequency is $\omega_{\mathrm{c}1}=2\pi\times 4.91$ GHz, with external decay rate $\kappa_{\mathrm{e}1}=2\pi\times 49.6$ kHz and intrinsic decay rate $\kappa_{\mathrm{i}1}=2\pi\times 298.9$ kHz. For the other microwave cavity (cavity-2), the intrinsic frequency is $\omega_{\mathrm{c}2}=2\pi\times 6.47$ GHz, accompanied by $\kappa_{\mathrm{e}2}=2\pi\times 21.7$ kHz and $\kappa_{\mathrm{i}2}=2\pi\times 245.7$ kHz. The oscillator is characterized to have frequency of $\omega_{\mathrm{m}}=2\pi\times 9.1$ MHz and the damping rate $\gamma_{\mathrm{m}}=2\pi\times 124$ Hz. Additionally, from the measurements based on frequency modulation technique, the single-photon coupling strengths are measured to be $g_{01}=2\pi\times 79$ Hz and $g_{02}=2\pi\times 46$ Hz. The detailed characterization method is described in Supplementary Note III.C.
@@
We choose the cavity-2 and the mechanical oscillator together to form the Floquet cavity system with modulation frequency $\omega_{\mathrm{m}}$, as schematically shown in Fig. 2c. It is worth noting that, in the absence of excitation, the mechanical oscillator is in a thermal state in equilibrium, with an extremely small time-averaged amplitude. To achieve Floquet modulation of the cavity-2 in a sinusoidal periodic manner, we designate the cavity-1 as auxiliary driving cavity to drive the mechanical oscillator into a self-induced oscillation or phonon lasing state (Supplementary Note IV) through optomechanical parametric excitation [55, 56, 44].
@@
We apply a blue-sideband drive at the frequency $\omega_{\mathrm{d}}=\omega_{\mathrm{c1}}+\omega_{\mathrm{m}}$ with drive power $P_{\mathrm{d}}$. The mechanical oscillator’s vibrations are gradually amplified and enter self-induced oscillation at a critical power $(P_{\mathrm{dc}})$ of $3\,\mathrm{dBm}$, as shown in Fig. 2d. Note that the power values correspond to the signal generator output powers and are reported in arbitrary units. We assume that the displacement is a classical state $x(t)=X_{\mathrm{m}}\cos(\omega_{\mathrm{m}}t+\phi_{\mathrm{m}})+X_{\mathrm{0}}$ [57]. Here, $X_{\mathrm{m}}$ is the classical amplitude and the $X_{\mathrm{0}}$ is the oscillation offset that is incorporated into $\omega_{\mathrm{c2}}$ (see Methods). Correspondingly, the modulation strength $A=\frac{g_{\mathrm{02}}X_{\mathrm{m}}}{x_{\mathrm{zpf}}}$ is controlled by the drive power $P_{\mathrm{d}}$.
@@
We characterize the Floquet sidebands of cavity-2 by using a vector network analyzer (VNA) to extract the reflection coefficient $S_{\mathrm{21}}=R=1-\sqrt{\kappa_{\mathrm{e}}}\frac{\langle a(t)\rangle}{\langle a_{\mathrm{in}}(t)\rangle}$. Combining with Eq. (1), the depth of the different sidebands is also determined by the Bessel function (detailed in Supplementary Note II.B), as shown in Fig. 2e. These sidebands are symmetrically distributed with a spacing corresponding to the oscillator frequency $\omega_{\mathrm{m}}$. We studied the effect of drive power, the results are shown in Fig. 2e for $P_{\mathrm{d}}$ at -5 dBm, 3 dBm, and 10 dBm. Before the critical power point $P_{\mathrm{dc}}$, the mechanical oscillator stays in thermal state and thus only the intrinsic resonance frequency of cavity-2 can be detected. When $P_{\mathrm{d}}>P_{\mathrm{dc}}$, the self-induced mechanical oscillation modulates cavity-2 such that it can be effectively described as the Floquet cavity. Increasing the modulation strength enhances the ability of photons to transition to higher-order Floquet energy levels. Consequently, the intensity of the intrinsic frequency signal diminishes, while higher-order sidebands become more pronounced. The result from the free parameter fitting indicates that, under a 10 dBm drive, $\beta$ is approximately 1.92.
@@
Figure 3: Emergence and evolution of Floquet comb. a Illustration of the setup. The blue arrow represents a fixed drive signal at $\Delta_{\mathrm{d}}=\omega_{\mathrm{m}}$ to excite mechanical oscillator. The red arrow in the Floquet cavity (cavity-2) represents the pump signal responsible for the generation of the frequency comb. b Emergence and evolution of frequency combs. Spectral responses are sorted into three types, and are marked in different colors as dots. We consider the frequency comb to be effective when the number of comb teeth $\geq 4$. The gray points represent the absence of the comb. The Floquet comb manifests in the regions marked with purple dots. In red-dotted region, cavity-2 itself produces a Kerr frequency comb. The black and red dashed lines are the dividing lines between the different regions. c Several examples of frequency comb corresponding to the dark points highlighted in b are shown, with each comb tooth marked by a line and a dot.

### v3

# Pump-Threshold-Free Frequency Comb via Cavity Floquet Engineering

@@ Abstract
Frequency combs have revolutionized communication, metrology, and spectroscopy. Considerable efforts have been devoted to developing integrated combs, primarily leveraging Pockels or Kerr nonlinearities. Here, we demonstrate an alternative frequency comb generated via cavity Floquet engineering. By periodically modulating the cavity resonance frequency through a driven mechanical oscillator, a Floquet cavity with multiple equally spaced frequency components is created. These sidebands exhibit nearest-neighbor coupling and are phase-locked to the external modulation drive. A pump tone interacts with the pre-modulated cavity to generate the output frequency comb, which we implement in an on-chip microwave cavity optomechanical system. This approach operates independently of a pumping threshold and is insensitive to pump detuning. Consequently, it enables comb generation under far-sideband pumping with nanowatt-scale total power consumption, providing an ultra-low-power platform for integrated frequency comb synthesis.
@@ Introduction
Frequency combs, composed of discrete, equally spaced frequencies [1, 2], have contributed to advancements in optical communication [3, 4], precision metrology [5, 6], spectroscopy [7, 8], and atomic clocks [9, 10, 11]. The compact on-chip frequency combs (microcombs) mainly consist of Kerr combs [12, 13, 14] and electro-optic (EO) combs [15, 16, 17, 18, 19]. Kerr combs relying on $\chi^{3}$ nonlinearity have been successfully demonstrated in high-$Q$ microresonators based on materials such as silica [12], silicon nitride [20], silicon carbide [21], diamond [22], and lithium niobate (LN) [23]. Integrated EO combs utilize the Pockels effect, which has also recently been realized in thin-film LN devices [15, 16, 17] and lithium tantalate [19]. Despite these remarkable advances, EO and Kerr combs are facing limitations of their generation efficiency [15, 24, 25] or tuning precision [26, 20]. Therefore, emerging paradigms for frequency comb generation are highly desired [27, 28, 29], as they offer promising avenues to overcome the aforementioned limitations in applications such as astrocombs [30, 31] and molecular fingerprinting [7, 32]. The frequency combs based on engineered nonlinearity could also offer a scalable and low-cost manner for emerging fields such as integrated photonic data links [33], quantum networks [34], and multiphoton entanglement [35, 36, 37]. Notably, Floquet engineering [38], which involves the periodic time-dependent modulation, has the capability to control and generate multi-mode sidebands by tuning energy spectra [39, 40, 41, 42, 43, 44, 45], as well as to induce interactions between modes [46, 47]. These equally spaced sidebands demonstrate that Floquet engineering is a promising, albeit underexplored, paradigm for frequency comb generation [48, 49], without necessitating Kerr nonlinearity or Pockels effect.
@@
In this work, we demonstrate a frequency comb generated via cavity Floquet engineering, termed the Floquet cavity frequency comb. Specifically, by periodically modulating the resonance frequency of a cavity, we realize a Floquet cavity with multiple equally spaced sidebands. These sidebands are coupled via nearest-neighbor interactions and can be further phase-locked to the external periodic modulation. Then, a single-frequency pump tone is added, interacting with this pre-modulated multi-mode structure to generate the output frequency comb. Additionally, the absolute positions of the comb teeth are referenced to the frequency of the pump tone, with the intensity approximately symmetrically distributed around the cavity’s intrinsic frequency. Furthermore, in contrast to Kerr or EO combs based on microresonators, which require precise frequency tuning to match the free spectral range, our approach is flexibly tunable, with the spacing and number of comb teeth determined by the modulation source. The frequency components of the Floquet cavity are internally predefined, enabling the frequency comb to operate without the need to surpass a pumping threshold.
Figure 1: Floquet cavity and frequency comb. a Examples of cavities with periodically modulated resonance frequencies. b Model of the Floquet cavity with the modulation frequency $\Omega$ and strength AA. c Time evolution of the intracavity field amplitude in the Floquet cavity obtained via numerical simulation. For demonstration purposes, cavity dissipation is neglected and $A\approx 25\Omega$. d The relative strength in the frequency domain of the Floquet cavity modes $\omega_{0}+m\Omega$ (where $m$ is the index of the Floquet states). The red dotted line marks the intrinsic frequency $\omega_{0}$. e Schematics of the spectral density of Floquet cavity quasi-energy with increasing modulation strength AA (where $n$ is the Fock basis). f The schematic diagram of the generation of Floquet cavity frequency comb. By injecting the pump tone into the pre-modulated cavity, photons are emitted across multiple modes, resulting in a frequency comb.
@@
Based on our theoretical framework, such a frequency comb can be realized across a wide range of physical systems. Here, we demonstrate its implementation for modulating the resonance frequency of a superconducting microwave cavity via a mechanically compliant capacitor. To achieve sinusoidal modulation, we additionally employ an auxiliary cavity to establish a conventional optomechanical coupling, driving the mechanical oscillator from thermal state into self-induced oscillation [50]. In contrast to Kerr optomechanical combs [51, 52, 53], which rely on the nonlinear coupling between phonons and photons [54, 55, 56], our approach achieves an extended tuning range for the pump signal, including frequencies that are substantially red-detuned from the cavity’s natural resonance. Furthermore, this approach reduces the pump power by a factor of approximately $10^{6}$, which is limited primarily by the detection efficiency. Cavity Floquet engineering enables efficient frequency comb generation even when the pump is far detuned from the cavity’s intrinsic frequency, with the total power reaching the nanowatt level. This thus provides an ultra-low-power platform for frequency comb generation.
@@ Results
We introduce a general Floquet cavity, defined as a cavity whose resonance frequency $\omega_{\mathrm{c}}$ is periodically modulated in time. It can be mapped to counterparts in well-established physical systems, including Fabry–Pérot resonators of a movable mirror or microwave circuits with adjustable capacitance or inductance, as illustrated in Fig. 1a. The resonance modes of these Floquet cavities transform from a single-frequency state to a superposition of multiple frequencies (see Supplementary Note I.A). The Hamiltonian of the Floquet cavity is given by $H_{\mathrm{Floquet}}=\mathrm{\hbar}\left(\omega_{\mathrm{0}}+A\cos(\Omega t+\phi_{0})\right)\hat{a}^{\dagger}\hat{a}$, where $\omega_{\mathrm{0}}$ is the intrinsic frequency, $A$ is the modulation strength, $\Omega$ is the modulation frequency, and we drop an irrelevant phase $\phi_{0}$, as illustrated in Fig. 1b.
In terms of Floquet theory, such periodic modulation of the cavity frequency effectively creates “Floquet sidebands” around the cavity’s intrinsic frequency. The quasi-energy of this Floquet cavity satisfies $\epsilon_{nm}/\hbar=n\omega_{\mathrm{0}}+m\Omega$, with the eigen-frequencies $\omega_{\mathrm{0}},\omega_{\mathrm{0}}\pm\Omega,\omega_{\mathrm{0}}\pm 2\Omega,\dots$ (see Methods), where integer $n$ is the Fock basis of bare cavity and $m$ is the Floquet index. We present a numerical example illustrating the time evolution of the intracavity field in the Floquet cavity in Fig. 1c. A set of equally spaced components corresponding to the intracavity field in the Fourier domain is shown in Fig. 1d. The relative strength of the Floquet modes is determined by $J_{m}(\beta)$, the Bessel function of the first kind with order $m$ (see Supplementary Note I.B and I.C). Here we define the dimensionless modulation coefficient $\beta=\frac{A}{\Omega}$. With a sufficiently large modulation strength, a substantial number of detectable Floquet modes can be produced.
Remarkably, since these modes are induced by Floquet engineering, they are not independent (as in modes in spatially separate cavities) but are correlated. The modulation strength $A/2$ determines the adjacent coupling strength between the different Floquet modes, thus forming what is referred to as a giant-mode cavity. A stronger modulation strength clearly enhances the generation of additional sidebands with appreciable intensity, as shown in Fig. 1e. Furthermore, this coupling enables energy input into the cavity to be transferred to different sidebands. It is notable that these sidebands inherently exhibit coherence, with their phases locked to the external modulation. Consequently, any input pump tone, even in the few-photon regime, is emitted across multiple sideband modes, forming a frequency comb with spacing $\Omega$, as illustrated in Fig. 1f.
Considering an input pump tone at the frequency $\omega_{\mathrm{in}}$, from the equation of motion, we obtain that the complex amplitude in the Floquet cavity satisfies (Supplementary Note II.A):
@@
$\times\sum_{m=-\infty}^{\infty}\sum_{l=-\infty}^{\infty}\frac{J_{l}(\beta)J_{m-l}(-\beta)}{il\Omega+\kappa/2-i\Delta}e^{im\Omega t},$
where $\Delta=\omega_{\text{in}}-\omega_{0}$ arises from the rotation of the frame, $\kappa$ represents the total dissipation of the cavity, $\kappa_{\mathrm{e}}$ denotes the external dissipation due to the coupling between the cavity and the external input/output, and $a_{\mathrm{in}}$ represents the externally injected pump field. For a coherent modulation source and pump tone, the output field consequently forms a set of isolated and coherent frequency-comb lines in the frequency domain. Noise from the pump and modulation sources causes linewidth broadening of the comb teeth, with details provided in Supplementary Notes II.B and II.C.
For a Floquet cavity frequency comb, the modulation coefficient $\beta$ is predefined and decoupled from $a_{\mathrm{in}}$, enabling frequency comb operation without a power threshold for the pumping tone, as indicated by Equation (1). In other terms, the pump tone does not induce nonlinear effects but instead serves as a probe for a system with a prescribed time-varying response function. Furthermore, we found that the pump tone exhibits flexible tuning capabilities within the approximate range $2\beta\Omega\lesssim\Delta\lesssim 2\beta\Omega$, as detailed in the Supplementary Note II.A. In contrast, in Kerr optomechanical combs, the generation of the frequency comb by optomechanical non-linearity depends on the pump tone itself, meaning $\beta$ is influenced by both $a_{\mathrm{in}}$ and $\Delta$. Consequently, $a_{\mathrm{in}}$ must reach a sufficient intensity to overcome an initiation threshold. Additionally, $\Delta$ needs to be constrained within a range that can effectively excite the oscillator (For $\kappa<\Omega$, $\Delta\approx\Omega$) in Kerr optomechanical combs. In terms of the comb spectral profile, EO and Kerr combs are typically symmetric about the pump frequency, whereas the Floquet cavity frequency comb is centered at the intrinsic cavity resonance ($\omega_{0}$), with a spectral envelope that is largely independent of the pump detuning.
@@
Figure 2: Construct the Floquet cavity in optomechanical systems. a Micrographs of the device, showing the entire chip and zoomed-in drums. We select the larger drum (left one) for the demonstration, while the other one is ignored. b Equivalent circuit model: straight arrows indicate the corresponding applied signals (blue for cavity-1, purple for cavity-2). c Cavity-1 serves as the auxiliary driving cavity to drive the mechanical oscillator into self-induced oscillation. Cavity-2 is modulated by the oscillator to construct the Floquet giant-mode cavity. d Mechanical oscillator spectrum transitioning from thermal state to self-induced oscillation regime; the red curve marks the spectrum at the critical drive power $P_{\mathrm{dc}}$. e The reflection spectrum of the cavity-2 measured by S21 as a function of increasing drive power. Colored dots: measurements; black lines: fits based on the reflection coefficient $R$. $X_{\mathrm{0}}$ is incorporated into $\omega_{\mathrm{c2}}$. The inset marks the driving frequency schematics.
We demonstrate a Floquet cavity frequency comb in a superconducting microwave circuit, as illustrated in Fig. 2a. The resonance frequency of the cavity is modulated by a displacement-dependent capacitor. The sample used in this study is similar to that reported in our previous work [57, 58] (detailed fabrication in Supplementary Note III.A). Our device consists of two aluminum drumhead mechanical oscillators [59, 60] optomechanically coupled to two microwave cavities, as schematically shown in Fig. 2b. Here, we ignore one of the mechanical oscillators. The Hamiltonian of such a system can be expressed as
@@
where $\omega_{\mathrm{c},j}$ represents the $j$-th cavity’s intrinsic frequency. $\omega_{\mathrm{m}}$ represents the mechanical resonant frequency, and $a_{j}$ ($a_{j}^{\dagger}$) and $b$ ($b^{\dagger}$) represent the annihilation (creation) operators of the cavities and mechanical oscillator, respectively. $\hat{x}$ is the displacement of the mechanical oscillator, and $x_{\mathrm{zpf}}$ represents the zero-point fluctuation of the mechanical oscillator’s displacement. Besides, the coefficient $g_{\mathrm{0},j}$ represents the single-photon coupling strength between the $j$-th cavity and mechanical oscillator.
Our entire device operates in a dilution refrigerator at a temperature of 10 mK (detailed setup in Supplementary Note III.B). The first cavity’s (cavity-1) intrinsic frequency is $\omega_{\mathrm{c}1}=2\pi\times 4.91~\text{GHz}$, with external decay rate $\kappa_{\mathrm{e}1}=2\pi\times 49.6~\text{kHz}$ and intrinsic decay rate $\kappa_{\mathrm{i}1}=2\pi\times 298.9~\text{kHz}$. For the other microwave cavity (cavity-2), the intrinsic frequency is $\omega_{\mathrm{c}2}=2\pi\times 6.47~\text{GHz}$, accompanied by $\kappa_{\mathrm{e}2}=2\pi\times 21.7~\text{kHz}$ and $\kappa_{\mathrm{i}2}=2\pi\times 245.7~\text{kHz}$. The oscillator is characterized to have frequency of $\omega_{\mathrm{m}}=2\pi\times 9.1~\text{MHz}$ and the damping rate $\gamma_{\mathrm{m}}=2\pi\times 124~\text{Hz}$. Additionally, from the measurements based on the frequency modulation technique, the single-photon coupling strengths are measured to be $g_{0,1}=2\pi\times 79~\text{Hz}$ and $g_{0,2}=2\pi\times 46~\text{Hz}$. The detailed characterization method is described in Supplementary Note III.C.
@@
We choose the cavity-2 and the mechanical oscillator together to form the Floquet cavity system with modulation frequency $\omega_{\mathrm{m}}$, as schematically shown in Fig. 2c. It is worth noting that, in the absence of excitation, the mechanical oscillator is in a thermal state in equilibrium, with an extremely small time-averaged amplitude and random phase. To achieve Floquet modulation of cavity-2 in a sinusoidal periodic manner, we designate cavity-1 as an auxiliary driving cavity to drive the mechanical oscillator into a self-induced oscillation (Supplementary Note IV) through optomechanical parametric excitation [50, 61, 62, 63].
@@
We apply a blue-sideband drive at the frequency $\omega_{\mathrm{d}}=\omega_{\mathrm{c1}}+\omega_{\mathrm{m}}$ with drive power $P_{\mathrm{d}}$. As the drive power increases, the mechanical oscillator’s vibrations are gradually amplified and enter self-induced oscillation at a critical power $(P_{\mathrm{dc}})$ of $3~\mathrm{dBm}$, as shown in Fig. 2d. Note that all reported power values correspond to the output power of the signal generator. The mechanical displacement of the self-induced oscillation can be described as a coherent oscillatory motion $x(t)=X_{\mathrm{m}}\cos(\omega_{\mathrm{m}}t+\phi_{\mathrm{m}})+X_{0}$. Here, $X_{\mathrm{m}}$ denotes a stable oscillation amplitude and $X_{0}$ is the equilibrium offset incorporated into the effective cavity resonance $\omega_{\mathrm{c2}}$ (see Methods). Correspondingly, the modulation strength $A=\frac{g_{02}X_{\mathrm{m}}}{x_{\mathrm{zpf}}}$ is governed by the external drive power $P_{\mathrm{d}}$. Furthermore, the oscillation phase exhibits coherence with the blue-sideband drive, arising from the underlying optomechanical phonon lasing mechanism [61, 62, 63, 64], which sustains stable periodic motion.
@@
We characterize the Floquet sidebands of cavity-2 by using a vector network analyzer to extract the reflection coefficient $S_{\mathrm{21}}=R=1-\sqrt{\kappa_{\mathrm{e}}}\frac{\langle a(t)\rangle}{\langle a_{\mathrm{in}}(t)\rangle}$. Combining with Equation (1), the depth of the different sidebands is also determined by the Bessel function, as shown in Fig. 2e. These sidebands are symmetrically distributed with a spacing corresponding to the oscillator frequency $\omega_{\mathrm{m}}$. We studied the effect of drive power, and the results are shown in Fig. 2e for $P_{\mathrm{d}}$ at 5-5 dBm, 3 dBm, and 10 dBm. Before the critical power point $P_{\mathrm{dc}}$, the mechanical oscillator stays in a thermal state, and thus only the intrinsic resonance frequency of cavity-2 can be detected. When $P_{\mathrm{d}}>P_{\mathrm{dc}}$, the self-induced mechanical oscillation, with stable amplitude and phase, modulates cavity-2 such that it can be effectively described as the Floquet cavity. Increasing the modulation strength enhances the ability of photons to transition to higher-order Floquet energy levels. Consequently, the intensity of the intrinsic frequency signal diminishes, while higher-order sidebands become more pronounced. The result from the free parameter fitting indicates that, under a 10 dBm drive, $\beta$ is approximately 1.92. Given the other device parameters, the corresponding mechanical oscillator amplitude $X_{\textrm{m}}$ is calculated as 1.2 nm.
Figure 3: Emergence and evolution of Floquet cavity frequency comb. a Illustration of the setup. The blue arrow represents a fixed drive signal at $\Delta_{\mathrm{d}}=\omega_{\mathrm{m}}$ to excite the mechanical oscillator. The red arrow in the Floquet cavity (cavity-2) represents the pump signal responsible for the generation of the frequency comb. b Emergence and evolution of frequency combs. Spectral responses are sorted into three types and are marked in different colors as dots. We consider the frequency comb to be effective when the number of comb teeth $\geq 4$. The gray points represent the absence of the comb. The Floquet comb manifests in the regions marked with purple dots. In the red-dotted region, cavity-2 itself produces a Kerr frequency comb. The black and red dashed lines are the dividing lines between the different regions. c Several examples of frequency comb corresponding to the dark points highlighted in b are shown, with each comb tooth marked by a line and a dot.